OpenStudy (anonymous):

Find the LCD: 5/ 7ab^5 , 2/ 7a^3b^3

OpenStudy (across):

You have that \(7ab^5=7\cdot a\cdot b\cdot b\cdot b\cdot b\cdot b\) and \(7a^3b^3=7\cdot a\cdot a\cdot a\cdot b\cdot b\cdot b\). Therefore, the LCD is \(7ab^3\).